OpenStudy (dimples97):

Write an equation in point-slope form of the line through point J(4, –4) with slope 4.

OpenStudy (amorfide):

standard equation of a line is y=mx+b where b is the y intercept, and m is the slope we know that the slope is 4 so m=4 y=4x+b standard coordinate is (x,y) we know the point (4,-4) so x=4, y=-4 substitute these values into y=4x+b and work out b

OpenStudy (mathmale):

This approach will certainly work! However, it'd be faster to begin with the general "point-slope form of the equation of a straight line and then substitute the given slope and point coordinates. \[y-y _{1}=m(x-x _{1})\]

OpenStudy (tkhunny):

Point-Slope: \((y-y_{0}) = m(x-x_{0})\) You are given the point, \((x_{0},y_{0})\) You are given the slope. Substitute the three values.

OpenStudy (mathmale):

Please begin that substitution now, and share your work.

OpenStudy (mathmale):

Would you please not just post a question and then log off. To be fair, stick around for a few minutes at least to see whether anyone tries to help you.

OpenStudy (sshayer):

eq. of a line is y=mx+b ...(1) if it passes through (x1,y1) y1=m x1+b ....(2) subtract (2) from (1) y-y1=m(x-x1)